# Multi-agent Adaptive Mechanism Design

### v1

@@ 4.3. Discussions
Truthfulness is weakly necessary for cost-optimality. Reducing cost of robustness requires accurate knowledge on the posterior distribution of peer’s observation $p^{*}(x_{j}\mid x_{i})$. Therefore, to ensure estimation is not biased, all players should be incentivized to stay truthful at all times. This provides yet another evidence on the necessity of truthfulness, in addition to Proposition 2.1.
@@
Robustness to fluctuation/non-stationarity of player performance. We assume each player’s skill (i.e., law $p_{i}(x_{i}\mid y)$) is consistent throughout the sequential tasks, an assumption not necessarily true in practice. Players may under or over-perform in certain rounds compared to other rounds, resulting in skill fluctuation and non-stationarity. In DRAM, we apply a distributionally robust mechanism each round. This robustness holds not only for estimation inaccuracy, but also to inaccuracy from other sources. This means as long as the actual reference distribution stays within the ambiguity set defined by $\hat{p}_{ik}$ and $\eta_{k}$, players are still incentivized to stay truthful. Indeed, the principal could even widen up or narrow the ambiguity set by adjusting $\eta_{k}$, looking for more robustness or less cost.
@@
Robustness to adversary. The distributionally robust mechanism would also provide robustness to adversarial behavior from players. When a player intentionally lies in a small portion of rounds, it would only slightly bias the estimation. As long as it does not surpass the ambiguity margin as designed in each epoch, the mechanism would not break down. In addition, the assignment procedure means an adversary would at most disrupt at most $2T$ out of $NT$ player–task interactions in total (being one focal player and one reference player), possibly spread across different players and tasks. Of course, here we only talk about the mechanism’s robustness, and an adversary would also contaminate the collected data, potentially biasing subsequent inference or learning procedures. For defending against adversaries in downstream tasks such as aggregation, there is a line of research that focuses on robust estimators and aggregators (gibbard1973manipulation; moulin1980strategy; caragiannis2016truthful; arieli2018robust).
@@
Flexibility with reference player assignment procedure. In DRAM, each player is assigned one corresponding reference player $j$, to which her reports will be compared. Any assignment procedure (deterministic or randomized) could be used for this process, and some could provide robustness to possible adverse players. As an example, suppose we use cyclic matching, where player $i+1$ is assigned to player $i$ as reference for $i<N$, and player $1$ is assigned to player $N$. Under cyclic matching, any adversary would disrupt at most two players, and the majority is not affected. Furthermore, at the beginning of each epoch, we could rerun the procedure and assign new players. Such replacement generates little extra computational costs, since we need to update the estimation and regenerate mechanism anyway.
@@
Compatibility with delayed/batched feedback. In the practical setting, the feedback to the principal might not be immediately available and may come in batches (chapelle2011empirical; mcmahan2013ad). The delayed/batched feedback setting has been studied in multiple online learning and decision-making problems (joulani2013online; gao2019batched). In DRAM, the mechanisms are computed at the beginning of each epoch, and stay the same throughout. This means DRAM naturally handles delayed and batched feedback since report data is only required for computation at the beginning of each epoch. Particularly, Corollary 4.3 suggests that $O(\log\log T)$ epochs are already sufficient for the $O(\sqrt{T})$ bound up to logarithmic terms. Nevertheless, such small epoch counts rely on a carefully designed epoch schedule. For example, DRAM uses a geometric epoch schedule, under which it quickly adapts early on, and slows down when sufficient data are gained. Deviation from the $O(\sqrt{T})$ bound may appear when the principal faces a different epoch schedule constraint (perchet2016batched).
@@ 5. Experiments
In this section, we perform numerical experiments to verify and demonstrate the effectiveness of our proposed algorithm.
Environment. We consider a sequential labeling game (as in Figure 1 and Example 3.1) with $N=3$ players and $d=3$ labels with a uniform prior $p_{Y}(y)=1/3$. Each player $i$ has a diagonally-dominant skill distribution $p_{i}(\cdot\mid y)$ that is symmetric across labels:
@@
and the player-wise robustness thresholds $\tilde{\eta}_{i}$ from Theorem 3.7, then set
@@
Given $\varepsilon=10^{-3}$, we plug $(\tilde{\eta}_{\text{used}},\rho_{\text{used}})$ into the warm-start length formula in Algorithm 1 to obtain $\tau$. For this setting, $\tau$ is on the order of $10^{5}$, so the warm-start phase occupies only a moderate fraction of the horizon. In the warm-start phase, we use the simple fact-checking mechanism: reward both player $1$ if their report agrees, and 0 if not.
@@
Truthfulness checks. We verify truthfulness via a retrospective approach. We set all participating players to be always truthful. At the beginning of every epoch, we perform a truthfulness check using the true joint distribution $p^{*}(X_{i},X_{j})$.
@@
We define the IC gap for player $i$ in epoch $k$ as
@@
and, for each episode, record the minimum $\mathrm{Gap}_{ik}$ across all players and epochs. If any $\mathrm{Gap}_{ik}$ is negative, we count the episode as an IC violation. Figure 2 shows the histogram of per-episode minimum IC gaps.
@@ Appendix A Deferred Proofs
In round $t$, denote player $i$’s report strategy by $s_{it}:\mathcal{Y}\to\Delta(\mathcal{Y})$, where $s_{it}(z\mid x)$ is the probability to report $z$ given $x$. Define player $i$’s information structure $\sigma_{it}$ as her conditional probability law to report $z$ conditional on true label $y$. Then we have
@@
Alternatively, we can write $\sigma=s_{it}\circ p_{i}$ as the information structure is induced by strategy $s_{it}$. We let $\boldsymbol{\sigma}=(\sigma_{1t},\dotsb,\sigma_{Nt})$ be the information structure profile of all players. Suppose the principal observes report $\mathbf{z}_{t}=\left(z_{1t},\cdots,z_{Nt-1}\right)$. The principal’s optimal expected accuracy under such report is
@@
since $A_{t}$ and $Y_{t}$ are independent conditional on $\boldsymbol{Z}_{t}$. Note that conditional law $\mathbb{P}\left(Y_{t}=y\mid\boldsymbol{Z}_{t}=\mathbf{z}\right)$ is decided by and only by players’ skills $p_{i}$ and report strategy $s_{it}$. Then the optimal expected accuracy under $\sigma$ is
@@
Note that this lemma applies for general $\boldsymbol{\sigma}$, even those where player $i$’s report may depend on other players’ observation. In our setting, we are only concerned with the subset of structures where each player’s report is independent from others’ observation.
@@
Necessity. Since we have a information structure profile $\boldsymbol{\sigma}$ that achieves maximal accuracy, from Lemma A.1 we know that for any player $i$, any information structure induced by a strategy $\sigma^{\prime}_{i}=s^{\prime}\circ p_{i}$ must be a garbling of $\sigma_{i}=s\circ p_{i}$. We now show that such property indicates that $s$ must be a permutation strategy $\pi$.
@@
Feasibility. Suppose $\mathbf{B}$ is invertible. Then for arbitrary matrix $\mathbf{M}$, there exists $\mathbf{R}=(\mathbf{B}^{-1}\mathbf{M})^{\intercal}$ such that $\mathbf{B}\mathbf{R}^{\intercal}=\mathbf{M}$. Notice that the entry $\mathbf{M}_{xy}$ is exactly player $1$’s expected reward given she observes label $x$ and reports label $y$. Hence for our purposes, we can construct an $\mathbf{M}$ whose diagonal entries are greater than $c$, and off-diagonal entries are less than $c$, then the corresponding $\mathbf{R}$ satisfies the first two constraints. (Actually, if $\mathbf{B}$ is invertible, the first two constraints of (2) can be satisfied for arbitrary right-hand side values.)
@@
To ensure player $i$ stays truthful, from Theorem 3.4 we need to find a margin $\delta$ such that $\delta/2\kappa^{*}\geq\eta$. We consider the best case $\delta/2\kappa^{*}=\eta$. Combining with Theorem 3.5 leads to
@@
Suppose player $i$’s observation $X_{it}=x$. Then from Bayes’ rule we have
@@
Estimation of the reference distribution $p^{*}(x_{j}\mid x_{i}).$ As suggested in Section 3, optimizing the cost of mechanism relies on accurate knowledge over the reference distribution $p^{*}(x_{j}\mid x_{i})$. Therefore, we first focus on accurate estimations on this distribution using players reports. Throughout this part, players’ reports are assumed to be truthful, i.e. $z_{i}=x_{i}$, so intuitively speaking principal should faithfully recover pp* if given enough data.
@@
Lemma A.2 (Concentration inequality of the empirical distribution (weissman2003inequalities)).
@@
Suppose the principal has received $T$ rounds of reports from player $i$ and $j$. Assuming players are always truthful. Let $\hat{p}(x_{j}\mid x_{i})$ be the empirical conditional distribution defined in Algorithm 1. Define the ambiguity set with ambiguity level $\eta$ as
@@
We first consider one label $x_{i}\in\mathcal{Y}$. Within $T$ rounds, the count of $x_{i}$ from player $i$’s report follows a binomial distribution. Let $\rho=\min_{x\in\mathcal{Y}}\mathbb{P}(X_{i}=x)$.
@@
Warm-starting. The ambiguity threshold $\tilde{\eta}$ is the smallest value across players, on the maximum ambiguity a distributionally robust mechanism can tolerate. For mathematical convenience we update the parameter to let $\tilde{\eta}<1/\sqrt{2}$, so that we have a cleaner bound in subsequent derivations. The fact-checking mechanism ensures truthfulness because of Lemma 4.1. The length of this phase is $O(\log(N\log T))$, which results in smaller order total regrets even when we have constant regret in each round.
Bounding the regret. Now we focus on the algorithm for a single player $i$. Consider what happens in epoch $k$, where $t\in(\tau_{k-1},\tau_{k}]$. At the beginning of the epoch, we have $\tau_{k-1}=2^{k}\tau$ data points, and we set $\eta_{k}$ in a specific way so that
@@
From Theorem 3.7, we know that the mechanism $R_{ik}$ acquired by solving Eq.(3) guarantees player $i$’s truthfulness when $p_{i}^{*}\in S_{\eta_{k}}(\hat{p}_{ik})$. Also, in the duration of epoch $k$, the expected regret for a single round is $c\cdot C_{1}\eta_{k}/(1-C_{2}\eta_{k})$ where $C_{1}$ and $C_{2}$ are constants as defined in Eq.(8). Therefore, the total expect regret for player $i$ across all $T$ rounds is
@@
Finally, we know that for one player in one epoch, the scheme ensures that applying union bound across all $N$ players and $\log T$ periods. Hence applying union bound, we know that truthfulness is held with probability $1-\varepsilon$. Also, note that the regret in the warm start phase is at most $O(\tau)$ and therefore dominated by the regret from the adaptive phase. Thus the total regret is $N$ times the single player’s regret and therefore $O(N\sqrt{T}\log(N(\log T)/\varepsilon)).$
@@
We start by considering the two-player, two-label case. Notice that as long as agents are perfectly truthful in the first $t$ rounds, on round $t$ the principal essentially has $t$ data points i.i.d. sampled from $p_{\boldsymbol{X}}$. In other words, the reward policy in previous rounds will not affect data distribution as long as it satisfies all the constraints. This specialty structure avoids cross-round (bandit-style) discussions on sequential policies. Instead, we can examine optimal policy round-by-round. In the following, we suppose the game is currently at round $T+1$, and all the previous rounds are truthful, so the principal has $T$ i.i.d. data points from $p_{\boldsymbol{X}}$. Without loss of generality we let cost $c=1$.
We first start by constructing a pair of problem instances $p_{0},p_{1}\in\Delta(\mathcal{X}^{N})$ that simultaneously satisfies the two requirements: i) the two instances are statistically close; ii) the optimal mechanisms differ sharply. Let the focal player be player $1$ and the reference player be $2$. We now study the reward mechanisms for the focal player $1$. Consider the following examples on $p_{0}$ and $p_{1}$, whereas from the focal player’s perspective, the two instances has:
@@
(Here the $1+2\delta$ is a relaxation on the bound $(a+d)/2<1+\delta$.) Notice that $\mathbf{M}_{xy}$ is exactly player $1$’s expected reward given she observes label $x$ and reports label $y$. Therefore, for any $\mathbf{M}_{0}$ that is cheap and satisfies the constraints under $p_{0}$, its performance under $p_{1}$ is $\mathbf{M}_{1}=\mathbf{B}_{1}\mathbf{B}_{0}^{-1}\mathbf{M}_{0}$. However, we have that:
@@
and we must have the one-player, single-round expected regret be greater than the same lower bound (since all mechanism with expected regret smaller than the lower bound would fail the test with probability greater than $\varepsilon.$)
@@
From Lemma 1 of (radanovic2013robust), it is known that for any mechanism with more than two players, its truthful Bayesian Nash Equilibrium can correspond to a 2-player mechanism with the same expected payment at its truthful Bayesian Nash Equilibrium. Therefore, more players would not bring additional benefits than the 2-player case, and hence we can simply apply the same lower bound for the N-player case. Hence the $N$-player mechanism design problem would have regret lower bound as:
@@
    Player $i$ does not observe $x_{i}$ for enough number of times. ($T_{x_{i}}$ is small).
@@
Suppose the principal has received $T$ rounds of reports from player $i$ and $j$. Assuming players are always truthful. Let $\hat{p}(x_{j}\mid x_{i})$ be the conditional distribution estimation from the general estimator in Definition 4.5. Let $\rho=\min_{x\in\mathcal{Y}}\mathbb{P}(X_{i}=x)$. Define the ambiguity set with ambiguity level $\eta_{\varepsilon}(\rho T/2)$ as
@@
Now we focus on the regret for a single player $i$. Consider what happens in epoch $k$, where $t\in(\tau_{k-1},\tau_{k}]$. At the beginning of the epoch, we have $\tau_{k-1}$ data points, and we set $\eta_{k}=\eta_{\varepsilon/Nm(d+1)}(\rho\tau_{k-1}/2)$. Thus we know from Lemma A.5 that $p_{i}^{*}\in S_{\eta_{k}}(\hat{p}_{ik})$ with probability at least $1-\varepsilon/(Nm)-(d+1)\exp(-\rho\tau_{k-1}/8)$.
From Theorem 3.7, we know that the mechanism $R_{ik}$ acquired by solving Eq.(3) guarantees player $i$’s truthfulness when $p_{i}^{*}\in S_{\eta_{k}}(\hat{p}_{ik})$. Also, in the duration of epoch $k$, the expected regret for a single round is $c\cdot C_{1}\eta_{k}/(1-C_{2}\eta_{k})$ where $C_{1}$ and $C_{2}$ are constants as defined in Eq.(8). Therefore, the total expect regret for player $i$ across all $T$ rounds is
@@
Finally, we know that for one player in one epoch, the scheme ensures that applying union bound across all $N$ players and $m$ periods. Hence applying union bound, we know that truthfulness is held with probability
@@
Also, note that the regret in the warm start phase is at most $O(\tau)$ and therefore dominated by the regret from the adaptive phase. Thus the total regret is $N$ times the single player’s regret and therefore

### v2

@@ 4.3. Discussions
Robustness to fluctuation/non-stationarity of agent performance. We assume each agent’s skill (i.e., law $p_{i}(x_{i}\mid y)$) is consistent throughout the sequential tasks, an assumption not necessarily true in practice. agents may under or over-perform in certain rounds compared to other rounds, resulting in skill fluctuation and non-stationarity. In DRAM, we apply a distributionally robust mechanism each round. This robustness holds not only for estimation inaccuracy, but also to inaccuracy from other sources. This means as long as the actual reference distribution stays within the ambiguity set defined by $\hat{p}_{ik}$ and $\eta_{k}$, agents are still incentivized to stay truthful. Indeed, the principal could even widen up or narrow the ambiguity set by adjusting $\eta_{k}$, looking for more robustness or less cost.
@@
Robustness to adversary. For the same reason, the distributionally robust mechanism would also provide robustness to adversarial behavior from agents. When an agent intentionally lies in a small portion of rounds, it would only slightly bias the estimation. As long as it does not surpass the ambiguity margin as designed in each epoch, the mechanism would not break down. In addition, the assignment procedure of reference agents may provide additional defense. An adversary would at most disrupt at most $2T$ out of $NT$ agent–task interactions in total (being one focal agent and one reference agent), possibly spread across different agents and tasks.
@@
Flexibility with reference agent assignment procedure. In DRAM, each agent is assigned one corresponding reference agent $j$, to which her reports will be compared. Any assignment procedure (deterministic or randomized) could be used for this process, and some could provide robustness to possible adverse agents. As an example, suppose we use cyclic matching, where agent $i+1$ is assigned to agent $i$ as reference for $i<N$, and agent $1$ is assigned to agent $N$. Under cyclic matching, any adversary would disrupt at most two agents, and the majority is not affected. Furthermore, at the beginning of each epoch, we could rerun the procedure and assign new agents. Such replacement generates little extra computational costs, since the principal needs to update estimation and regenerate mechanism anyway.
@@
Compatibility with delayed/batched feedback. In the practical setting, the feedback to the principal might not be immediately available and may come in batches (Chapelle and Li, 2011; McMahan et al., 2013). The delayed/batched feedback setting has been studied in multiple online learning and decision-making problems (Joulani et al., 2013; Gao et al., 2019). In DRAM, the mechanisms are computed at the beginning of each epoch, and stay the same throughout. This means DRAM naturally handles delayed and batched feedback since report data is only required for computation at the beginning of each epoch. Particularly, Corollary 4.3 suggests that $O(\log\log T)$ epochs are already sufficient for the $O(\sqrt{T})$ bound up to logarithmic terms. Nevertheless, such small epoch counts rely on a carefully designed epoch schedule. For example, DRAM uses a geometric epoch schedule, under which it quickly adapts early on, and slows down when sufficient data are gained. Deviation from the $O(\sqrt{T})$ bound may appear when the principal faces a different epoch schedule constraint (Perchet et al., 2016).
@@ 5. Experiments
In this section, we perform numerical experiments to verify and demonstrate our proposed algorithm.
Environment. We consider a sequential labeling game (as in Figure 1 and Example 3.1) with $N=3$ agents and $d=3$ labels with a uniform prior $p_{Y}(y)=1/3$. Each agent $i$ has a diagonally-dominant skill distribution $p_{i}(\cdot\mid y)$ that is symmetric across labels:
@@
and the agent-wise robustness thresholds $\tilde{\eta}_{i}$ from Theorem 3.7, then set
@@
Given $\varepsilon=10^{-3}$, we plug $(\tilde{\eta}_{\text{used}},\rho_{\text{used}})$ into the warm-start length formula in Algorithm 1 to obtain $\tau$. For this setting, $\tau$ is on the order of $10^{5}$, so the warm-start phase occupies only a moderate fraction of the horizon. In the warm-start phase, we use the simple fact-checking mechanism: reward both agent $1$ if their report agrees, and 0 if not.
@@
Truthfulness checks. We verify truthfulness via a retrospective approach. We set all participating agents to be always truthful. At the beginning of every epoch, we perform a truthfulness check using the true joint distribution $p^{*}(X_{i},X_{j})$.
@@
We define the IC gap for agent $i$ in epoch $k$ as
@@
and, for each episode, record the minimum $\mathrm{Gap}_{ik}$ across all agents and epochs. If any $\mathrm{Gap}_{ik}$ is negative, we count the episode as an IC violation. Figure 2 shows the histogram of per-episode minimum IC gaps.
@@ Appendix A Deferred Proofs
In round $t$, denote agent $i$’s report strategy by $s_{it}:\mathcal{Y}\to\Delta(\mathcal{Y})$, where $s_{it}(z\mid x)$ is the probability to report $z$ given $x$. Define agent $i$’s information structure $\sigma_{it}$ as her conditional probability law to report $z$ conditional on true label $y$. Then we have
@@
Alternatively, we can write $\sigma=s_{it}\circ p_{i}$ as the information structure is induced by strategy $s_{it}$. We let $\boldsymbol{\sigma}=(\sigma_{1t},\dotsb,\sigma_{Nt})$ be the information structure profile of all agents. Suppose the principal observes report $\mathbf{z}_{t}=\left(z_{1t},\cdots,z_{Nt-1}\right)$. The principal’s optimal expected accuracy under such report is
@@
since $A_{t}$ and $Y_{t}$ are independent conditional on $\boldsymbol{Z}_{t}$. Note that conditional law $\mathbb{P}\left(Y_{t}=y\mid\boldsymbol{Z}_{t}=\mathbf{z}\right)$ is decided by and only by agents’ skills $p_{i}$ and report strategy $s_{it}$. Then the optimal expected accuracy under $\sigma$ is
@@
Note that this lemma applies for general $\boldsymbol{\sigma}$, even those where agent $i$’s report may depend on other agents’ observation. In our setting, we are only concerned with the subset of structures where each agent’s report is independent from others’ observation.
@@
Necessity. Since we have a information structure profile $\boldsymbol{\sigma}$ that achieves maximal accuracy, from Lemma A.1 we know that for any agent $i$, any information structure induced by a strategy $\sigma^{\prime}_{i}=s^{\prime}\circ p_{i}$ must be a garbling of $\sigma_{i}=s\circ p_{i}$. We now show that such property indicates that $s$ must be a permutation strategy $\pi$.
@@
Feasibility. Suppose $\mathbf{B}$ is invertible. Then for arbitrary matrix $\mathbf{M}$, there exists $\mathbf{R}=(\mathbf{B}^{-1}\mathbf{M})^{\intercal}$ such that $\mathbf{B}\mathbf{R}^{\intercal}=\mathbf{M}$. Notice that the entry $\mathbf{M}_{xy}$ is exactly agent $1$’s expected reward given she observes label $x$ and reports label $y$. Hence for our purposes, we can construct an $\mathbf{M}$ whose diagonal entries are greater than $c$, and off-diagonal entries are less than $c$, then the corresponding $\mathbf{R}$ satisfies the first two constraints. (Actually, if $\mathbf{B}$ is invertible, the first two constraints of (2) can be satisfied for arbitrary right-hand side values.)
@@
To ensure agent $i$ stays truthful, from Theorem 3.4 we need to find a margin $\delta$ such that $\delta/2\kappa^{*}\geq\eta$. We consider the best case $\delta/2\kappa^{*}=\eta$. Combining with Theorem 3.5 leads to
@@
Suppose agent $i$’s observation $X_{it}=x$. Then from Bayes’ rule we have
@@
Estimation of the reference distribution $p^{*}(x_{j}\mid x_{i}).$ As suggested in Section 3, optimizing the cost of mechanism relies on accurate knowledge over the reference distribution $p^{*}(x_{j}\mid x_{i})$. Therefore, we first focus on accurate estimations on this distribution using agents reports. Throughout this part, agents’ reports are assumed to be truthful, i.e. $z_{i}=x_{i}$, so intuitively speaking principal should faithfully recover pp* if given enough data.
@@
Lemma A.2 (Concentration inequality of the empirical distribution (Weissman et al., 2003)).
@@
Suppose the principal has received $T$ rounds of reports from agent $i$ and $j$. Assuming agents are always truthful. Let $\hat{p}(x_{j}\mid x_{i})$ be the empirical conditional distribution defined in Algorithm 1. Define the ambiguity set with ambiguity level $\eta$ as
@@
We first consider one label $x_{i}\in\mathcal{Y}$. Within $T$ rounds, the count of $x_{i}$ from agent $i$’s report follows a binomial distribution. Let $\rho=\min_{x\in\mathcal{Y}}\mathbb{P}(X_{i}=x)$.
@@
Warm-starting. The ambiguity threshold $\tilde{\eta}$ is the smallest value across agents, on the maximum ambiguity a distributionally robust mechanism can tolerate. For mathematical convenience we update the parameter to let $\tilde{\eta}<1/\sqrt{2}$, so that we have a cleaner bound in subsequent derivations. The fact-checking mechanism ensures truthfulness because of Lemma 4.1. The length of this phase is $O(\log(N\log T))$, which results in smaller order total regrets even when we have constant regret in each round.
Bounding the regret. Now we focus on the algorithm for a single agent $i$. Consider what happens in epoch $k$, where $t\in(\tau_{k-1},\tau_{k}]$. At the beginning of the epoch, we have $\tau_{k-1}=2^{k}\tau$ data points, and we set $\eta_{k}$ in a specific way so that
@@
From Theorem 3.7, we know that the mechanism $R_{ik}$ acquired by solving Eq.(LABEL:eq:_dist-robust_two-agent_mechanism_design) guarantees agent $i$’s truthfulness when $p_{i}^{*}\in S_{\eta_{k}}(\hat{p}_{ik})$. Also, in the duration of epoch $k$, the expected regret for a single round is $c\cdot C_{1}\eta_{k}/(1-C_{2}\eta_{k})$ where $C_{1}$ and $C_{2}$ are constants as defined in Eq.(7). Therefore, the total expect regret for agent $i$ across all $T$ rounds is
@@
Finally, we know that for one agent in one epoch, the scheme ensures that applying union bound across all $N$ agents and $\log T$ periods. Hence applying union bound, we know that truthfulness is held with probability $1-\varepsilon$. Also, note that the regret in the warm start phase is at most $O(\tau)$ and therefore dominated by the regret from the adaptive phase. Thus the total regret is $N$ times the single agent’s regret and therefore $O(N\sqrt{T}\log(N(\log T)/\varepsilon)).$
@@
We start by considering the two-agent, two-label case. Notice that as long as agents are perfectly truthful in the first $t$ rounds, on round $t$ the principal essentially has $t$ data points i.i.d. sampled from $p_{\boldsymbol{X}}$. In other words, the reward policy in previous rounds will not affect data distribution as long as it satisfies all the constraints. This specialty structure avoids cross-round (bandit-style) discussions on sequential policies. Instead, we can examine optimal policy round-by-round. In the following, we suppose the game is currently at round $T+1$, and all the previous rounds are truthful, so the principal has $T$ i.i.d. data points from $p_{\boldsymbol{X}}$. Without loss of generality we let cost $c=1$.
We first start by constructing a pair of problem instances $p_{0},p_{1}\in\Delta(\mathcal{X}^{N})$ that simultaneously satisfies the two requirements: i) the two instances are statistically close; ii) the optimal mechanisms differ sharply. Let the focal agent be agent $1$ and the reference agent be $2$. We now study the reward mechanisms for the focal agent $1$. Consider the following examples on $p_{0}$ and $p_{1}$, whereas from the focal agent’s perspective, the two instances has:
@@
(Here the $1+2\delta$ is a relaxation on the bound $(a+d)/2<1+\delta$.) Notice that $\mathbf{M}_{xy}$ is exactly agent $1$’s expected reward given she observes label $x$ and reports label $y$. Therefore, for any $\mathbf{M}_{0}$ that is cheap and satisfies the constraints under $p_{0}$, its performance under $p_{1}$ is $\mathbf{M}_{1}=\mathbf{B}_{1}\mathbf{B}_{0}^{-1}\mathbf{M}_{0}$. However, we have that:
@@
and we must have the one-agent, single-round expected regret be greater than the same lower bound (since all mechanism with expected regret smaller than the lower bound would fail the test with probability greater than $\varepsilon.$)
@@
From Lemma 1 of (Radanovic and Faltings, 2013), it is known that for any mechanism with more than two agents, its truthful Bayesian Nash Equilibrium can correspond to a 2-agent mechanism with the same expected payment at its truthful Bayesian Nash Equilibrium. Therefore, more agents would not bring additional benefits than the 2-agent case, and hence we can simply apply the same lower bound for the N-agent case. Hence the $N$-agent mechanism design problem would have regret lower bound as:
@@
    agent $i$ does not observe $x_{i}$ for enough number of times. ($T_{x_{i}}$ is small).
@@
Suppose the principal has received $T$ rounds of reports from agent $i$ and $j$. Assuming agents are always truthful. Let $\hat{p}(x_{j}\mid x_{i})$ be the conditional distribution estimation from the general estimator in Definition 4.5. Let $\rho=\min_{x\in\mathcal{Y}}\mathbb{P}(X_{i}=x)$. Define the ambiguity set with ambiguity level $\eta_{\varepsilon}(\rho T/2)$ as
@@
Now we focus on the regret for a single agent $i$. Consider what happens in epoch $k$, where $t\in(\tau_{k-1},\tau_{k}]$. At the beginning of the epoch, we have $\tau_{k-1}$ data points, and we set $\eta_{k}=\eta_{\varepsilon/Nm(d+1)}(\rho\tau_{k-1}/2)$. Thus we know from Lemma A.5 that $p_{i}^{*}\in S_{\eta_{k}}(\hat{p}_{ik})$ with probability at least $1-\varepsilon/(Nm)-(d+1)\exp(-\rho\tau_{k-1}/8)$.
From Theorem 3.7, we know that the mechanism $R_{ik}$ acquired by solving Eq.(LABEL:eq:_dist-robust_two-agent_mechanism_design) guarantees agent $i$’s truthfulness when $p_{i}^{*}\in S_{\eta_{k}}(\hat{p}_{ik})$. Also, in the duration of epoch $k$, the expected regret for a single round is $c\cdot C_{1}\eta_{k}/(1-C_{2}\eta_{k})$ where $C_{1}$ and $C_{2}$ are constants as defined in Eq.(7). Therefore, the total expect regret for agent $i$ across all $T$ rounds is
@@
Finally, we know that for one agent in one epoch, the scheme ensures that applying union bound across all $N$ agents and $m$ periods. Hence applying union bound, we know that truthfulness is held with probability
@@
Also, note that the regret in the warm start phase is at most $O(\tau)$ and therefore dominated by the regret from the adaptive phase. Thus the total regret is $N$ times the single agent’s regret and therefore